OpenStudy (anonymous):

What happens to the area of a parallelogram when you double both the height and the base? The area is halved The area is doubled The area is quadrupled The area is cut in half

OpenStudy (anonymous):

@MissSmartiez @welshfella

OpenStudy (anonymous):

The answer is B

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

sure!

OpenStudy (welshfella):

try an example base 2 and height 3 - changes to base 4 and height 6

OpenStudy (welshfella):

Its not B

OpenStudy (anonymous):

yes it is

OpenStudy (welshfella):

work it out for yourself

OpenStudy (welshfella):

2 *3 = 6 4 8 6 = 24

OpenStudy (anonymous):

Who do I listen to?

OpenStudy (welshfella):

* 4 * 6 = 24

OpenStudy (anonymous):

A parallelogram is basiclly 2 triangles which would be a parrallelogram and if it is doubled well basiclly its either B or C

OpenStudy (welshfella):

area = base * height 2 * 3 = 6 4 * 6 = 24 what is the relation between 6 and 24?

OpenStudy (anonymous):

6 goes into 24? @welshfella

OpenStudy (welshfella):

6 * what = 24?

OpenStudy (anonymous):

4 @welshfella

OpenStudy (welshfella):

right so is that doubled or quadrupled?

OpenStudy (anonymous):

quadrupled? @welshfella

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

that is what I said a while ago

OpenStudy (anonymous):

I said B or C

OpenStudy (welshfella):

the area will be 4 times greater - quadrupled

OpenStudy (anonymous):

Thank you @welshfella

OpenStudy (welshfella):

yw